STUDIES OF THE PREDISSOCIATED, QUASILINEAR B^{1}A′ STATE OF CH^{35}Cl AND CD^{35}Cl BY OPTICAL-OPTICAL DOUBLE RESONANCE SPECTROSCOPY

C. TAO, C. MUKARAKATE AND \underline{S. A. REID}, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

Last year at this meeting, we reported studies of the predissociated, quasilinear B^{1}A′ state of fluorocarbene (CHF and CDF) using a fluorescence dip detected optical-optical double resonance technique via the A^{1}A′′ state.^{a} Recently, we have extended these observations to chlorocarbene, measuring OODR spectra of both CH^{35}Cl and CD^{35}Cl. By recording OODR spectra of both isotopomers, we are able to pinpoint the origin of the B^{1}A′ state, while lies near 22 400 cm^{−1}. In contrast to CHF, the B^{1}A′ origin in chlorocarbene lies below the energetic threshold of the lowest dissociation channel, C(^{3}P) + HCl. However, at higher energies several channels open up, including C-H and C-Cl bond fission, and the OODR spectra show significant lifetime broadening. Trends in the measured linewidths with energy will be discussed.

aC. Tao, S. A. Reid, T. W. Schmidt, and S. H. Kable, J. Chem. Phys. 125, 051105 (2007).